# DLBacktrace: A Model Agnostic Explainability for any Deep Learning Models

### v1

@@ 3.1 Introduction
(a) Sample Network
(b) Relevance Ouptput for Sample Network
Figure 1: Illustration Depecting Backtrace Calculation for a Sample Network
Backtrace is a technique for analyzing neural networks that involves tracing the relevance of each component from the output back to the input.
This approach clarifies how each element contributes to the final prediction. By distributing relevance scores across various layers, Backtrace provides insights into feature importance, information flow, and potential biases, which facilitates improved model interpretation and validation without relying on external dependencies.
Backtrace has the following advantages over other available tools:
  • No dependence on a sample selection algorithm :
  • No dependence on a secondary white-box algorithm :
    The relevance is calculated directly from the network itself. This prevents any variation in importance due to type, hyperparameters and assumptions of secondary algorithms.
  • Deterministic in nature
    The relevance scores won’t change on repeated calculations on the same sample. Hence, can be used in live environments or training workflows as a result of its independence from external factors.
3.2 Methodology
Backtrace operates in two modes: Default Mode and Contrast Mode.
First we describe the Basic Methodology of Backtrace in Default Mode as follows :
3.2.1 Basic Methodology
Every neural network consists of multiple layers. Each layer has a variation of the following basic operation:
$y=\Phi(Wx+b)$
where,
  • $\Phi$ = activation function
  • $W$ = weight matrix of the layer
  • $b$ = bias
  • $x$ = input
  • $y$ = output
This can be further organized as:
$y=\Phi(X_{p}+X_{n}+b)$
where,
  • $X_{p}=\sum W_{i}x_{i}\quad\forall W_{i}x_{i}>0$
  • $X_{n}=\sum W_{i}x_{i}\quad\forall W_{i}x_{i}<0$
Activation functions can be categorized into monotonic and non-monotonic functions.
  • Non-Monotonic functions: The relevance is propagated as is.
  • Monotonic functions: The relevance is switched off for positive or negative components based on saturation.
3.2.2 Relevance Propagation
The aforementioned modes represent the basic operations at each source layer for propagating relevance to the destination layer. The procedure for relevance calculation is as follows:
@@ 3.2.2 Relevance Propagation
Note: Any loss of relevance during propagation is due to network bias.
The relevance of a single sample represents local importance. For global importance, the relevance of each feature can be aggregated after normalization at the sample level.
@@ 3.3 Algorithm
The algorithm has two modes of operation:
  • Default Mode
  • Contrastive Mode
3.3.1 Default Mode
In this mode, a single relevance is associated with each unit. The relevance is propagated by proportionately distributing it between positive and negative components. If the relevance associated with $y$ is $r_{y}$ and with $x$ is $r_{x}$, then for the $j$th unit in $y$, we compute:
$T_{j}=X_{pj}+|X_{nj}|+|b_{j}|$ (1)
@@ 3.3.1 Default Mode
$R_{pj}$ and $R_{nj}$ are distributed among $x$ in the following manner:
@@ 3.3.2 Contrastive Mode:
In this mode, each unit is assigned dual relevance, distributed between positive and negative components. This approach facilitates separate analyses of supporting and detracting influences. Unlike single-mode propagation, which combines relevance into aggregated scores, dual-mode propagation provides clarity by isolating favorable and adverse contributions. This separation enhances interpretability, enabling deeper insights into features that negatively impact predictions a capability essential for identifying counterfactuals or assessing model biases in high-stakes scenarios.
If the relevance associated with $y$ are $r_{yp},r_{yn}$ and with $x$ are $r_{xp},r_{xn}$, then for the $j$th unit in $y$, we compute:
@@
We then calculate Determine $R_{pj}$, $R_{nj}$, and Relevance Polarity as described in Algorithm 1.
@@ 3.4 Relevance for Attention Layers:
Currently, the majority of AI models across various applications are primarily based on the attention mechanism [36]. Accordingly, we have extended our Backtrace algorithm to provide support for this attention model [37].
Attention mechanism allows the model to focus on specific parts of the input sequence, dynamically weighting the importance of different elements when making predictions. The attention function employs the equation:
@@
For Multi-Head Attention, the implementation is as follows:
$MultiHead(Q,K,V)=\textbf{Concat}\left(head_{1},head_{2},...,head_{n}\right){W^{O}}$ (8)
where each head is computed as
@@
such that
  • Q, K, V: Query, Key, Value Matrices
  • $W_{i}^{Q}$, $W_{i}^{K}$, $W_{i}^{V}$: Weight matrices for the i-th head
  • $W^{O}$: Weight matrix for combining all the heads after concatenation
  • Concat: Concatenation of the outputs from all attention heads
3.4.1 Relevance Propagation for Attention Layers
Suppose the input to the attention layer is $x$ and the output is $y$. The relevance associated with $y$ is $r_{y}$. To compute the relevance using the Backtrace, we use the steps as indicated in Algorithm 3 below:
@@ 4 Benchmarking
In this section, we present a comparative study to benchmark our proposed Backtrace algorithm against various existing explainability methods. The goal of this evaluation is to assess the effectiveness, robustness, and interpretability of Backtrace in providing meaningful insights into model predictions across different data modalities, including tabular, image, and text data. By systematically comparing our approach with established methods, we aim to highlight the advantages and potential limitations of Backtrace in the context of explainable artificial intelligence (XAI).
@@ 4.2 Metrics
4.2.1 Tabular Modality
@@ 4.2.3 Text Modality
To evaluate the textual modality, we use the Token Perturbation for Explanation Quality (ToPEQ) metric, which assesses the robustness of model explanations by analyzing the impact of token perturbations. We employ the Least Relevant First AUC (LeRF AUC) and Most Relevant First AUC (MoRF AUC) to measure sensitivity to the least and most important tokens, respectively. Additionally, we calculate Delta AUC, the difference between LeRF AUC and MoRF AUC, to further indicate the model’s ability to distinguish between important and unimportant features.
@@
4.3 Experiments
4.3.1 Tabular Modality
We evaluated 1,024 samples from the test set of the Lending Club dataset, using a fine-tuned MLP checkpoint that attained an accuracy of 0.89 and a weighted average F1 score of 0.87. We assessed Backtrace against widely used metrics for tabular data, specifically LIME and SHAP [4], and employed MPRT for comparison, along with Complexity to examine the simplicity of the model explanations as illustrated in Appendix A.1.1.
Table 1: Explaination Performance metrics for explanation methods - LIME,SHAP and Backtrace, including Mean values and feature contributions across different layers (fc1 to fc4). Lower values in MPRT and Complexity indicate better performance.
@@ 4.3.1 Tabular Modality
Backtrace 0.562 0.579 0.561 0.557 0.552 2.201
The proposed method, Backtrace, as demonstrated in Table 1, achieves superior performance compared to both LIME and SHAP, evidenced by lower Maximal Perturbation Robustness Test (MPRT) values across various layers. This highlights its improved interpretability and robustness in explainability. However, Backtrace exhibits higher computational complexity, reflecting the fine-grained, higher entropy of its explanations. This trade-off suggests the need to balance the quality of interpretability with the simplicity of model explanations.
4.3.2 Image Modality
We conducted an evaluation on 500 samples from the CIFAR-10 test set using a supervised, fine-tuned ResNet-34 model, which achieved a test accuracy of 75.85%. We compared Backtrace against several methods as illustrated in Appendix A.1.2, including Grad-CAM, vanilla gradient, smooth gradient, and integrated gradient. The comparison utilized metrics such as Faithfulness Correlation, Max Sensitivity, and Pixel Flipping.
@@ 4.3.2 Image Modality
Backtrace 0.199 0.617 0.199
The Evaluations conducted as shown in Table 2 reveals that Backtrace significantly outperforms traditional methods like Grad-CAM, Vanilla Gradient, Smooth Gradient, and Integrated Gradient across all key metrics. Backtrace achieves a superior Faithfulness Correlation score of (0.199), indicating a stronger alignment between its explanations and the model’s behavior. Additionally, it demonstrates robust performance with much lower Max Sensitivity (0.617) and Pixel Flipping (0.199) scores, highlighting its stability against input perturbations and better robustness in preserving the model’s predictive integrity under pixel modifications. Overall, Backtrace establishes itself as a more reliable and robust explainability technique for image modality tasks.
4.3.3 Text Modality
For the text modality, evaluation was performed on the evaluation set of the SST-2 dataset using a fine-tuned BERT model [^1] Model checkpoint used https://huggingface.co/textattack/bert-base-uncased-SST-2, achieving an F1 score of 0.926. Since explainable AI (XAI) for BERT and other transformer-based models is relatively new, we employed metrics based on token perturbation for explanation quality, specifically LeRF, MoRF, and Delta AUC, as introduced in [37]. We used methods like IG, SmoothGrad, AttnRoll, GradCAM and Input Grad as illustrated in Appendix A.1.5.
@@ 4.3.3 Text Modality
As shown in Table 3, Integrated Gradients (IG) delivered the strongest performance, achieving the lowest MoRF AUC and the highest LeRF and Delta AUC, underscoring the robustness of its explanations and precise feature attribution. Smooth Grad and Attn Roll also demonstrated commendable performance. Backtrace exhibited balanced results across LeRF and Delta AUC metrics, with a MoRF AUC of 15.431, showcasing its ability to provide meaningful explanations while allowing scope for further enhancement. These findings highlight the effectiveness of IG in explainability for transformer-based models while recognizing Backtrace’s potential as a promising and competitive approach.
4.4 Observations
@@ 5 Discussion
Additional illustrations of Backtrace for various use cases are provided in Appendix A.
5.1 Advantages of Backtrace
@@ 5.1.2 Feature Importance
  • We can also evaluate the dependence within the input source as the weight assignment happens on a per unit basis.
@@ 5.2 Applicability
The Backtrace framework is applicable in the following use-cases:
@@ 5.2.3 Fairness and bias analysis using the feature-wise importance
This is in continuation of the global importance of features. Based on the global importance of sensitive features (e.g. gender, age, etc.) and their alignment with the data, it can be inferred whether the model or data have undue bias towards any feature value.
@@ 5.2.5 Validating the model outcome
During deployment, validation of outcomes is extremely important for complete autonomous systems. The layer-wise relevance can be used for accomplishing this. The relevance for each layer is mapped in the vector space of the same dimension as the layer outcome, yet it is linearly related to the model outcome.
Since the information changes as it passes through the network, the relevance from lower layers, even input layers, can be used to get different outcomes. These outcomes can be used to validate the model outcome. The layers are generally multi-dimensional, for which either proximity-based methods or white-box regression algorithms can be used to derive outcomes.
@@ 7 Future Works
Future research on DLBacktrace will aim to broaden its use and improve how it scores relevance. Key areas for development include adapting DLBacktrace for complex and evolving model architectures, like advanced transformers and multimodal models to ensure it remains effective across different AI applications. Additionally, we aim to reduce the inference time making DLBacktrace more suitable for real-time applications in production environments such as autonomous systems and dynamic decision-making scenarios.
Future development will also explore the use of DLBacktrace for specific model improvements, including diagnosis and targeted editing. For example, DLBacktrace could assist in model pruning by identifying less critical components, tehreby optimizing model performance and efficiency. In addition, DLBacktrace’s potential for targeted model improvements will be explored. It can assist in model pruning, especially for Mixtures of Experts (MoEs), by identifying underutilized components or redundant experts to optimize performance and efficiency. It can also help in facilitate model merging, providing insights for seamless integration of multiple models, and layer swapping, enabling selective replacement of layers to enhance adaptability or performance. We also plan to apply DLBacktrace to out-of-distribution (OOD) detection, where it can help distinguish instances that fall outside the model’s training data, enhancing the robustness and reliability of AI systems.
Furthermore, extending DLBacktrace’s support for model-agnostic explainability will allow it to be seamlessly applied across various architectures, making it a versatile tool in explainable AI. These improvements will make DLBacktrace more useful and establish it as an important tool for understanding and improving models across a wide range of AI applications and tasks.
@@ A.1.2 Image Modality : Multi-Class Classification
(a) Original Image
(b) Backtrace
(c) Smooth Grad
(d) Vanilla Gradient
(e) IG
(f) GradCAM

### v2

@@ 3.1 Introduction
DLBacktrace is a technique for analyzing neural networks that involves tracing the relevance of each component from the output back to the input revealing the feature importance of input, the importance of each layer, information flow, and biases. It operates independently of external algorithms or datasets, ensuring consistency and deterministic results. Relevance is computed directly from the network, making DLBacktrace ideal for live environments and training workflows, with improved interpretation and validation. Utilizing DLBacktrace enables the extraction of local relevance for each sample, highlighting the importance unique to that specific instance. By aggregating the normalized scores across multiple samples, it effectively captures the global relevance. The procedure for relevance calculation for a given input sample and model (with trained weights) as illustrated in Figure  2 is as follows:
@@
3.2 DLBacktrace Modes
DLBacktrace works in two modes: Default and Contrastive. In Default Mode, a single relevance score is distributed proportionally across positive and negative components, influenced by saturation effects in activation functions. Contrastive Mode assigns dual relevance scores (positive and negative) to each unit, separating supportive and detractive contributions. This mode facilitates deeper analysis of negatively impactful features and enables bias detection or counterfactual evaluation in applications like sentiment analysis.
@@ 3.3 Algorithm
Context: Every deep learning model consists of multiple layers. Each layer has a variation of the basic operation:
$y=\Phi(Wx+b)$
where, $\Phi$ = activation function , $W$ = weight matrix of the layer , $b$ = bias , $x$ = input , $y$ = output .
This can be further organized as:
$y=\Phi(X_{p}+X_{n}+b)$
where, $X_{p}=\sum W_{i}x_{i}\quad\forall W_{i}x_{i}>0$ and $X_{n}=\sum W_{i}x_{i}\quad\forall W_{i}x_{i}<0$
3.3.1 Default Mode :
[…]
$T_{j}=X_{pj}+X_{nj}+b_{j}$ (1)
@@ 3.3.1 Default Mode :
$R_{pj}$ and $R_{nj}$ are distributed among $x$ as follows:
@@ 3.3.2 Contrastive Mode :
[…]
@@
We then determine $R_{pj}$, $R_{nj}$, and Relevance Polarity as described in Algorithm 1.
@@ 3.4 Relevance for Attention Layers:
Most modern AI models rely on the attention mechanism [29], which has been extended in our work to support the DLBacktrace algorithm.The attention mechanism allows the model to focus on relevant parts of the input sequence by dynamically adjusting the importance of elements for predictions. The attention function is defined as:
@@
For Multi-Head Attention, the function is:
$MultiHead(Q,K,V)=\textbf{Concat}\left(head_{1},\ldots,head_{n}\right)W^{O}$ (8)
@@
Where, $Q,K,V$: Query, Key, and Value matrices, $W_{i}^{Q},W_{i}^{K},W_{i}^{V}$: Weight matrices for each head,$W^{O}$: Weight matrix for combining all heads after concatenation and Concat: Concatenates the outputs from all heads .
3.4.1 Algorithm
Suppose the input to the attention layer is $x$ and the output is $y$. The relevance associated with $y$ is $r_{y}$. To compute the relevance using the DLBacktrace, we use the steps as indicated in Algorithm 3 below:
@@ 3.4.1 Algorithm
Note: Any loss of relevance during propagation is due to network bias.
4 Usage of DLBacktrace Library
DLBacktrace supports both PyTorch and TensorFlow frameworks. It can be easily installed using pip. Overall Workflow is as illustrated in Figure LABEL:fig:pipeline-fig.
1pip install dl-backtrace
For TensorFlow models, we can import and initialize the library as follows : (assume we have a pre-trained deep learning model)
1from dl_backtrace.tf_backtrace import DLBacktrace as B
2backtrace = B(model=model)
Similarly, for Pytorch-based models :
1from dl_backtrace.pytorch_backtrace import DLBacktrace as B
2backtrace = B(model=model)
To compute layer-wise outputs and evaluate relevance as illustrated in Figure 2 :
1layer_outputs = backtrace.predict(test_data[0])
2relevance = backtrace.eval(layer_outputs, mode=’default’, scaler=1, thresholding=0.5, task="binary-classification")
The relevance evaluation process in DLBacktrace involves adjusting four key attributes: mode, scaler, thresholding, model-type, and task.
The mode attribute defines the method for evaluating relevance (e.g., "default," "contrastive"), determining how layer contributions are measured. The model-type is used for specific model types that require special operations during graph creation (e.g., ’BERT’).
The scaler attribute controls the scaling of relevance scores; it provides initial relevance value at the last layer (e.g., scaler=1).
The thresholding attribute sets a cutoff for model confidence values, discarding predictions to select a predicted class below a specified threshold (e.g., thresholding=0.5 it is mostly used for segmentation task).
The task attribute specifies the type of task being performed (e.g., "binary-classification", "regression", "binary-segmentation"), influencing how relevance is computed based on the model task.
@@ 5 Benchmarking
In this section, we present a comparative study to benchmark our proposed DLBacktrace algorithm against various existing explainability methods. The goal of this evaluation is to assess the effectiveness, robustness, and interpretability of DLBacktrace in providing meaningful insights into model predictions across different data modalities, including tabular, image, and text data. By systematically comparing our approach with established methods, we aim to highlight the advantages and potential limitations of DLBacktrace in the context of explainable artificial intelligence (XAI).
@@ 5.2 Metrics
(a) DLBacktrace
(b) LIME
(c) SHAP
5.2.1 Tabular Modality:
For Tabular Modality we use MPRT and Complexity.
@@ 5.2.1 Tabular Modality:
(a) Original Image
(b) Backtrace
(c) Smooth Grad
(d) Vanilla Gradient
(e) IG
(f) GradCAM
@@ 5.2.2 Image Modality
For Image Modality we use Faithfulness Correlation, Max-Sensitivity and Pixel Flipping .
@@ 5.2.3 Text Modality
To evaluate the textual modality, we use the Token Perturbation for Explanation Quality (ToPEQ) metric, which assesses the robustness of model explanations by analyzing the impact of token perturbations.
@@
Table 1: Explaination Performance metrics for explanation methods - LIME,SHAP and DLBacktrace, including Mean values and feature contributions across different layers (fc1 to fc4). Lower values in MPRT and Complexity indicate better performance.
@@
DLBacktrace 0.562 0.579 0.561 0.557 0.552 2.201
5.3 Experiments
5.3.1 Tabular Modality
We evaluated 1,024 samples from the test set of the Lending Club dataset, using a fine-tuned MLP checkpoint that attained an accuracy of 0.89 and a weighted average F1 score of 0.87. We assessed DLBacktrace against widely used metrics for tabular data, specifically LIME and SHAP [11], and employed MPRT for comparison, along with Complexity to examine the simplicity of the model explanations as illustrated in Fig 3.
The proposed method, DLBacktrace, as demonstrated in Table 1, achieves superior performance compared to both LIME and SHAP, evidenced by lower Maximal Perturbation Robustness Test (MPRT) values across various layers. This highlights its improved interpretability and robustness in explainability. However, DLBacktrace exhibits higher computational complexity, reflecting the fine-grained, higher entropy of its explanations. This trade-off suggests the need to balance the quality of interpretability with the simplicity of model explanations.
@@ 5.3.1 Tabular Modality
DLBacktrace 0.199 0.617 0.199
@@
5.3.2 Image Modality
We conducted an evaluation on 500 samples from the CIFAR-10 test set using a supervised, fine-tuned ResNet-34 model, which achieved a test accuracy of 75.85%. We compared DLBacktrace against several methods as illustrated in Fig. 4, including Grad-CAM, vanilla gradient, smooth gradient, and integrated gradient. The comparison utilized metrics such as Faithfulness Correlation, Max Sensitivity, and Pixel Flipping.
The Evaluations conducted as shown in Table 2 reveals that DLBacktrace significantly outperforms traditional methods like Grad-CAM, Vanilla Gradient, Smooth Gradient, and Integrated Gradient across all key metrics. DLBacktrace achieves a superior Faithfulness Correlation score of (0.199), indicating a stronger alignment between its explanations and the model’s behavior. Additionally, it demonstrates robust performance with much lower Max Sensitivity (0.617) and Pixel Flipping (0.199) scores, highlighting its stability against input perturbations and better robustness in preserving the model’s predictive integrity under pixel modifications. Overall, DLBacktrace establishes itself as a more reliable and robust explainability technique for image modality tasks.
5.3.3 Text Modality
For the text modality, evaluation was performed on the evaluation set of the SST-2 dataset using a fine-tuned BERT model, achieving an F1 score of 0.926. Since explainable AI (XAI) for BERT and other transformer-based models is relatively new, we employed metrics based on token perturbation for explanation quality, specifically LeRF, MoRF, and Delta AUC, as introduced in [32]. We used methods like IG, SmoothGrad, AttnRoll, GradCAM and Input Grad as illustrated in Fig. 13.
As shown in Table 3, Integrated Gradients (IG) delivered the strongest performance, achieving the lowest MoRF AUC and the highest LeRF and Delta AUC, underscoring the robustness of its explanations and precise feature attribution. Smooth Grad and Attn Roll also demonstrated commendable performance. DLBacktrace exhibited balanced results across LeRF and Delta AUC metrics, with a MoRF AUC of 15.431, showcasing its ability to provide meaningful explanations while allowing scope for further enhancement. These findings highlight the effectiveness of IG in explainability for transformer-based models while recognizing DLBacktrace’s potential as a promising and competitive approach.
5.3.4 Observations
@@ 6 Discussion
DLBacktrace framework offers key advantages in model interpretation. It improves network analysis by addressing the limitations of traditional methods like distribution graphs and heatmaps, which only show node activations without reflecting their impact on the final prediction. It also distinguishes between input data and internal network biases. For feature importance, DLBacktrace assigns weights to input sources, making it easier to understand their influence on predictions, and avoids the issues seen with methods like Integrated Gradients and Shapley values. Additionally, it incorporates uncertainty by evaluating node weight distributions relative to past predictions.
DLBacktrace is versatile and applicable in several areas. It provides local and global feature importance, with global importance averaged across samples. In network analysis, it offers insights into layer relevance, including bias-to-input ratios and activation saturation, guiding network adjustments. The framework also supports fairness and bias analysis by evaluating the influence of sensitive features like gender or age. It helps assess process compliance by comparing feature rankings with business rules or regulations.
6.1 DLBacktrace Advantages
DLBacktrace has the following advantages over other available tools:
  • No dependence on a sample selection algorithm :
  • No dependence on a secondary white-box algorithm :
    The relevance is calculated directly from the network itself. This prevents any variation in importance due to type, hyperparameters, and assumptions of secondary algorithms.
  • Deterministic in nature
    The relevance scores won’t change on repeated calculations on the same sample. Hence, it can be used in live environments or training workflows as a result of its independence from external factors.
@@ 6.1.2 Feature Importance
  • We can also evaluate the dependence within the input source as the weight assignment happens on a per-unit basis.
@@ 6.2 Applicability
The DLBacktrace framework is applicable in the following use cases:
@@ 6.2.3 Fairness and bias analysis using the feature-wise importance
This is in continuation of the global importance of features. Based on the global importance of sensitive features (e.g., gender, age, etc.) and their alignment with the data, it can be inferred whether the model or data has an undue bias towards any feature value.
@@ 6.2.5 Validating the model outcome
During deployment, validation of outcomes is extremely important for complete autonomous systems. The layer-wise relevance can be used to accomplish this. The relevance for each layer is mapped in the vector space of the same dimension as the layer outcome, yet it is linearly related to the model outcome.
Since the information changes as it passes through the network, the relevance from lower layers, even input layers, can be used to get different outcomes. These outcomes can be used to validate the model outcome. The layers are generally multi-dimensional, for which proximity-based methods or white-box regression algorithms can derive outcomes.
@@ 8 Future Works
Future research on DLBacktrace will focus on expanding its applicability and improving its relevance scoring. Key areas include adapting it for complex models like advanced transformers and multimodal systems, ensuring its effectiveness across different AI applications. Efforts will also be made to reduce inference time for real-time use in production environments such as autonomous systems. We will explore it’s usage for model optimization, including pruning by identifying non-essential components. Additionally, we will also explore it’s usage for out-of-distribution (OOD) detection to enhance model robustness.
9 Ethical Concerns
DLBacktrace has the potential to support compliance with regulatory requirements, such as the EU GDPR Article 22, which grants individuals the right to an explanation regarding automated decision-making processes that significantly affect them. DLBacktrace helps ensure that AI systems remain accountable and meet these legal standards by providing transparent insights into model decisions. To foster responsible use, it is essential that DLBacktrace is implemented with the appropriate safeguards, including clear communication with users about how their data is being utilized and efforts to prevent any misuse. This study exclusively uses publicly available datasets and tools from open domains, adhering to their respective licensing and usage terms. No personal or sensitive data is involved in the study. Furthermore, the codebase used for this research is already publicly accessible, promoting transparency and reproducibility.